OpenStudy (anonymous):

A giant scoop, operated by a crane, is in the shape of a hemisphere of radius 21 inches. The scoop is filled with molten steel. The steel is then poured into a cylindrical storage tank with diameter 28 inches. How high will the molten steel rise?

OpenStudy (anonymous):

volume of hemisphere = 2/3 π r^3 = 2/3 *22/7 *21 *21 *21 = 19404 inch^3

OpenStudy (radar):

You now have the volume of molten steel, now calculate how high it will be when all poured into that cylinder.

OpenStudy (anonymous):

i got 19396

OpenStudy (anonymous):

volume of cylinder = π r^2h Here volume has to be equal to the volume of the hemispherical scoop so 19404 inch^3 = 22/7 * 28 cm * 28 cm * h => 19404 inch^3 = 2464 cm^2 * h => 19404 inch^3/2464 inch^2 = h => 7.88 inch = h So the molten steel will rise 7.88 inch high

OpenStudy (anonymous):

@Hikari you can check my calculations with a computer and find them correct. U might have used π as 3.14.....

OpenStudy (anonymous):

pls click GOOD ANSWER button ( :))